Larry has $20,000 in student loans. Part of this amount is at 4% simple interest the rest is at 6% simple interest. At the end of the first year the interest chargers were $920. Find the amount Larry borrowed at each rate. (Note: Larry did not make any payments on his loan the first year.)
Answer by mananth(16949)

let x be @6%
y @ 4%
..
x+y =20000......................1
..
0.06x+0.04y=920
multiply by 100
6x+4y=92000......................2
multiply (1) by -6
-6x-6y=-120000
add this to (2)
-6x+6x-6y+4y=-120000+92000
-2y=-28,000
/-2
y=14,000 amount @ 4%
balance $6000 @ 6%
...
m.ananth@hotmail.ca